OpenStudy (anonymous):

I need help solving this algebra problem. Shade the solution set to the system of inequalities. X + 2y > -2 2x + y<= 5

OpenStudy (anonymous):

I think the question is asking you to draw the graph for both equation then shade the solution.

OpenStudy (anonymous):

I have 4 graphs that each have shading for my answer to choose from. I'm hoping that if I can work the problem out it will lead me to which graph to choose.

OpenStudy (calculusfunctions):

OK, first draw the graphs of x + 2y = -2 and 2x + y = 5. Use a dotted line for the first equation because it doesn't have an equal sign.

OpenStudy (calculusfunctions):

Alright, then do see choices with those lines, where only the line for x + 2y = -2 is dotted?

OpenStudy (anonymous):

Let me look closely at each option.

OpenStudy (anonymous):

Putting each equation in to slope-intercept form (y=mx+b) first helps.

OpenStudy (anonymous):

All 4 choices show that. I'm trying to understand exactly which portion of the graph should be shaded. All lines, solid & dotted are set the same on each graph.

OpenStudy (anonymous):

One is a strictly greater than, so that must be a dotted line. The other is a ≤ so is a solid line. When you have them in y=mx+b form, the inequality signs imply shading in the y-direction, viz. vertical. > & ≥ is above the line, and < & ≤ is below the line.

OpenStudy (anonymous):

Ok. I think it is the graph with the shading in top right (almost) half corner. I only have 3 problems left to do & my brain is shutting down on me. I thank you for your help. You may see me again in a min!! ;)